OpenStudy (anonymous):

x2 + 4x + 8 = 0 This has to do with solving quadratic equations with complex numbers.

OpenStudy (anonymous):

then what do you do?

OpenStudy (anonymous):

do you have to solve it?

OpenStudy (anonymous):

like factoring?

OpenStudy (anonymous):

You can either complete the square, or you can use the quadratic formula which will lead you to your complex roots.

OpenStudy (anonymous):

\[x^2 + 4x + 8 = 0 \]\[x^2+4x=-8\] \[(x+2)^2=-8+4=-4\] \[x+2=\pm \sqrt{-4}=\pm2i\] \[x=-2\pm2i\]